# ProTrix: Building Models for Planning and Reasoning over Tables with Sentence Context

### v2

@@ Abstract
Tables play a crucial role in conveying information in various domains. We propose a Plan-then-Reason framework to answer different types of user queries over tables with sentence context. The framework first plans the reasoning paths over the context, then assigns each step to program-based or textual reasoning to reach the final answer. This framework enhances the table reasoning abilities for both in-context learning and fine-tuning methods. GPT-3.5-Turbo following Plan-then-Reason framework surpasses other prompting baselines without self-consistency while using less API calls and in-context demonstrations. We also construct an instruction tuning set TrixInstruct to evaluate the effectiveness of fine-tuning with this framework. We present ProTrix model family by finetuning models on TrixInstruct. Our experiments show that ProTrix family generalizes to diverse unseen tabular tasks with only 6k training instances. We further demonstrate that ProTrix can generate accurate and faithful explanations to answer complex free-form questions. Our work underscores the importance of the planning and reasoning abilities towards a model over tabular tasks with generalizability and interpretability. We open-source our dataset and models at https://github.com/WilliamZR/ProTrix.
@@ 1 Introduction
Tables, serve as a fundamental tool for organizing and presenting information across various domains. Whether in business reports, or scientific publications, tables are commonly employed to convey complex data effectively. Despite their widespread utility, the process of human beings answering questions involving tables appears to be time-consuming, given the often substantial amount of content involved. Recognizing this challenge, there arises a need to leverage Large Language Models (LLMs) to understand and respond to user queries automatically.
Figure 1: Demonstrations of user queries to a table in Wikipedia. The table is extracted from the Wikipedia page Athletics at the 2001 Goodwill Games. Some of the sentences with hyperlinks to the table are presented as sentence context.
@@
The last query also requires the model to recall commonsense knowledge to decide which contextual information can be used as evidence to judge if Kenya is successful at the 2001 Goodwill Games. Subsequently, the model must generate explanations to arrive at certain conclusions. The first two queries mainly require the model to fill the information gap in the query with a short-form answer while the third query seeks for information on a more general concept. The queries underscore the importance of planning and reasoning abilities to connect the general concepts with the actual information in the context and generate faithful and accurate explanations for conclusions.
@@
There are generally two ways to enhance reasoning ability. One is textual reasoning which prompts the model to answer questions step-by-step Wei et al. (2022). The other one is program-based reasoning, prompting the model to write code to answer the questions Chen et al. (2022). Each of the reasoning methods has its disadvantages as shown in Figure 2. The textual reasoning method such as Chain-of-Thought Wei et al. (2022) can be used to enhance the tabular reasoning ability but often lacks precision in tabular operations such as sorting, counting and filtering as shown in the right example in Figure 2, and may not generalize well to large tables Chen (2023). The program-based reasoning method can reason with high precision with SQL or Python code Chen et al. (2022). The left example in Figure 2 queries the color comparison between steaks with different cooking methods which is program-unsolvable. Therefore, it raises a need to integrate the advantages of program-based and textual reasoning.
In this paper, we propose a Plan-then-Reason framework for tabular tasks with sentence context that plans upon various types of user queries and then reasons over context to reach the final answers. The model could write code to extract necessary information from the table or perform specific operations with high precision, which would help the model generalize to unseen or larger tables. The model could leverage textual reasoning to maintain an understanding of general concepts and combine information from table and sentence context to reach final answers. Our experiments show that Plan-then-Reason surpasses existing prompting methods without self-consistency and requires less in-context examples and API calls.
@@
Meanwhile, few finetuning methods are designed for enhancing both planning and reasoning abilities of open-source models while these abilities are crucial for building a tabular model with generalizability and interpretability. Various pre-trained models are proposed for tabular tasks Yin et al. (2020); Wang et al. (2021); Iida et al. (2021); Deng et al. (2022); Yang et al. (2022); Jiang et al. (2022); Liu et al. (2021). But they are often limited to specific query types and could not generalize well to unseen tasks. Regarding models fine-tuned with respect to general tabular querying tasks Xie et al. (2022); Liu et al. (2023a); Zhang et al. (2023a); Zhuang et al. (2024), they are expected to generate the answers directly, which inevitably lacks interpretability. TableLLM assigns program-based or textual reasoning based on table size to answer the queries without planning based on the query and context Zhang et al. (2024).
@@
Recent base models are pre-trained with a large amount of corpora thus obtaining intrinsic commonsense knowledge Touvron et al. (2023); Roziere et al. (2023). These models suit as the backbones for our models that can fill the gap between queries and tables, understand general concepts, and plan the reasoning paths over table and sentence context. We construct an instruction tuning dataset TrixInstruct based on benchmarks with queries that are program-unsolvable or need combining information from table and sentence context. We finetune Llama-2-7B Touvron et al. (2023) and CodeLlama-7B Roziere et al. (2023) with TrixInstruct. The resulting models, the ProTrix [^1] Protrix originally means a chemical reactor for small-scale production with compatibility and process control. family is designed to Plan and Reason On TabulaR tasks with integration of code executIon and teXtual reasoning. Our experiments show that models trained with Plan-then-Reason framework can generalize to unseen tabular tasks in different domains with only a handful of training examples and give accurate and faithful explanations even for complex how and why questions.
@@
$\bullet$ We propose a Plan-then-Reason framework to answer user queries on tabular tasks with sentence context. The framework first plans the reasoning pathways by ingesting the query and the context, and assigns each step to textual and program-based reasoning to arrive at the final answer. We experiment with GPT-3.5-turbo to evaluate the effectiveness of this framework and find it surpasses existing methods without self-consistency.
@@ 2.1 Problem Formulation
This study centers on tabular tasks with sentence context. Each instance is structured as $(Q,T,S,A)$, where $Q$ represents users’ query. $T$ denotes a singular table, while $S$ denotes the sentence context. The sentence context is usually passages linked to the table or retrieved from a knowledge base. Finally $A$ stands for the predicted answer. The answer could be short-form for answering questions like hom many… or is this true or false…. For how and why questions, the answer is generally one or more sentences which is defined as free-form answers.
Figure 3: Illustration of our framework. The figure shows the process of our framework to answer a question. The framework first performs strategical planning, decomposing questions into reasoning chains to be solved by either table or sentence context, then perform reasoning based on symbolic programs or natural language to reach answers.
@@ 2.2 Plan-then-Reason
We propose a Plan-then-Reason framework for answering queries over tables and texts. The framework first ingests the query and the context by recalling common knowledge and general concepts. Then it begins to design the model’s reasoning pathway, planning the utilization of program-based and textual reasoning to arrive at conclusions.
@@
Then the model adaptively plans the reasoning path with program-based and textual reasoning to address the limitations of each reasoning method. For the first query, the model plans to use SQL to extract relevant cells from the table and make comparisons of concepts through textual reasoning. For the second query, the model decomposes the task into a multi-hop reasoning chain. It uses SQL to extract the silver medalist from the table and uses sentence context to verify his Olympic record.
@@
The reasoning phase initiates with program-based reasoning, writing SQL queries to extract relevant cells or perform operations such as counting and sorting. After running SQL on a code interpreter, the results are fed back into the model’s ongoing reasoning process. Subsequently, during textual reasoning, the model selects relevant sentences from noisy context to complement the table context. Reuben Kosgei… youngest ever winner of an Olympic gold medal suggests that he has won a gold medal in his career. At last, the model summarizes insights from the program-based and textual reasoning to reach the answer.
@@ 2.3 In-Context Learning
One intuitive way to evaluate the effectiveness of Plan-then-Reason work is through in-context learning. We annotate the planning and reasoning process and prompt LLMs to follow the pattern. We experiment with GPT-3.5-turbo on WikiTQ Pasupat and Liang (2015), FEVEROUS Aly et al. (2021) and TabFact Chen et al. (2019) following prompt shown in Table 1011 and 12 in Appendix.
@@ 2.4 Instruction Tuning
Based on the analysis in §1, we highlight the abilities Plan-then-Reason framework possesses towards tabular tasks with sentence context. (1) Understanding user’s query: use parametric knowledge of commonsense insights and general concepts to analyze the relationship between the query and the context; (2) Adaptive planning: decompose a query into sub-questions and plan to answer these sub-questions with different types of context or design multi-hop reasoning paths for the query, and (3) Blending program-based and textual reasoning: synthesize the strength of each method to maximize reasoning performance.
To train our model with such abilities, we construct an instruction tuning set TrixInstruct based on two benchmarks i.e., WikiTQ Pasupat and Liang (2015) and FEVEROUS Aly et al. (2021). WikiTQ involves table question-answering based on a single table, requiring multi-step reasoning and performing diverse data operations such as comparison, aggregation, and arithmetic computation. WikiTQ also contains cases that can not be solved by SQL programming solely Shi et al. (2020) which need to be solved with textual reasoning as the left example in Figure 2. Finetuning with such cases equips our models with the ability to plan between textual or program-based reasoning.
@@
For each task, we sample 4,000 instances from the training datasets. We employ GPT-4 Achiam et al. (2023) to generate responses according to the Plan-then-Reason framework following the prompts in Table 10 and 11 in Appendix. We filter out instances that GPT-4 predicts answers inconsistent with the original annotations. This results in a training set comprising 3,157 cases from FEVEROUS and 2,866 cases from WikiTQ. We train our models based on Llama-2-7B Touvron et al. (2023) and CodeLlama-7B Roziere et al. (2023).
@@ 3.1 Benchmarks for Evaluation
WikiSQL and WikiTQ are question answering benchmarks on tables from Wikipedia without sentence context Zhong et al. (2017); Pasupat and Liang (2015). HybridQA Chen et al. (2020) requires models to answer questions based on both tables and sentences. We use retrieved sentences, admittedly noisy, from Chen et al. (2020) as the sentence context. TATQA Zhu et al. (2021) is focused on tables with sentence context from financial reports.
@@
We follow our method in §2.4 to construct the evaluation dataset for FEVEROUS Aly et al. (2021). TabFact Chen et al. (2019) verifies claims based on tables from the Wikipedia. SciTAB Lu et al. (2023) focuses on tables from scientific papers. This benchmark requires compositional reasoning and commonsense knowledge.
@@
FetaQA contains what questions with multiple answers and how and why questions that requires model to generate explanations Nan et al. (2022). The original FetaQA dataset has highlighted relevant cells, we turn to a more challenging and realistic scenario where the highlighted cells are not provided as input and the model will answer the question directly based on the complete table context. Since our model is only finetuned on short-form answer tasks, FetaQA can be utilized to further evaluate the interpretability and generalizability of our models.
@@ 3.2 Short-form Answer Tasks
We choose the following baselines: (1) Closed-source model: We use the end-to-end QA performance of GPT-3.5-turbo and GPT-4 as baselines and compare our results on GPT-3.5-turbo with Binder Cheng et al. (2022b), ReAcTable Zhang et al. (2023b), StructGPT Jiang et al. (2023), Dater Ye et al. (2023), Chain-of-Table Wang et al. (2024) and TabSQLify Nahid and Rafiei (2024). (2) Finetuned SOTA: Please refer to Appendix A for details of finetuned SOTA methods (3) 7B parameter models: We first compare our model with the zero-shot performance of base models, Llama-2-7B Touvron et al. (2023) and CodeLlama 7B Roziere et al. (2023). Then we compare our model with TableLlama Zhang et al. (2023a) and TableLLM Zhang et al. (2024), which are the most similar baselines to our model and share the same base models [^2] We do not use StructLM Zhuang et al. (2024) as baseline since it is finetuned on most of our evaluation benchmarks while we want to compare with existing models in terms of generalizability.. TableLlama is a generalist model for end-to-end QA. TableLLM originally chooses textual or program-based reasoning for different benchmarks. To test the generalizability for different queries, we choose the reasoning method of TableLLM by table size. We use textual reasoning for tables less than 500 tokens and program-based reasoning for larger tables. We provide a detailed comparison with other closed-source models (Codex, PaLM2) and open-source model (Llama-3-8B) in Table 9 in Appendix.
@@
For fact verification, we match the predicted veracity label in the concluding sentence. And we use the evaluator from Cheng et al. (2022b) to evaluate the performance of Plan-then-Reason with GPT-3.5-turbo on WikiTQ. For 7B parameter models, we find that the heuristic matching evaluation for question-answering tasks would lead to overestimation or underestimation since our model is not finetuned to follow the grammar of gold answers in each dataset. Therefore, we deploy Llama-2-70B-chat Touvron et al. (2023) to predict if the concluding sentence answers the question correctly similarly as Zhang et al. (2024) following prompt in Table 13 in Appendix. Our evaluation results using LLM are checked manually and they align better with human evaluation of the concluding sentence. We report three-class F1 score for SciTAB and accuracy for other datasets.
@@
The result with GPT-3.5-turbo [^3] We use GPT-3.5-turo-16k-0613 for experiment. We only experiment with WikiTQ, TabFact and FEVEROUS due to limited budgets. in Table 1 suggests that Plan-then-Reason framework is effective for answering different types of queries over tables. Our method surpasses all the existing baselines without self-consistency on WikiTQ. It also obtains accuracy higher than Dater uses 20 responses and Binder uses 50 responses for self-consistency. Our method still falls behind Mix SC which predicts answers based on 5 responses with textual reasoning and 5 with program-based reasoning. But the self-consistency mechanism leads to a much larger computing cost. Our framework also surpasses existing baselines on TabFact except for StructGPT, demonstrating its effectiveness for fact verification tasks. Plan-then-Reason can also generalize to tables with sentence context. It increases the accuracy on FEVEROUS by 4.8%.
@@
The experimental result with Llama-2-7B and CodeLlama-7B in Table 1 shows that our finetuned models generalize to diverse tabular tasks with only 6k training instances. Compared with the backbone model Llama-2-7B, the performance gain of ProTrix on in-domain benchmarks is 34.8% on WikiTQ and 28.7% on FEVEROUS. And the performance gain on out-of-domain benchmarks is 21.5% on average. Comparing the out-of-domain performance with TableLlama, ProTrix surpasses TableLlama by 25.7% on WikiSQL, 15.8% on SciTAB, 9.6% on HybridQA and 11.8% on TATQA. The overall performance gain on out-of-domain benchmarks demonstrates the planning and reasoning abilities obtained from TrixInstruct is not restricted to in-domain tasks. Our finetuned model ProTrix adaptatively generalizes to queries with different input and output configurations and can even be applied to specific domains such as science and finance.
ProTrix-Coder still falls behind TableLLM on question answering task with sentence context since TableLLM is finetuned with 8k cases from TATQA. We are surprised to find that ProTrix-Coder surpasses TableLLM on WikiTQ and WikiSQL by 4.9% and 7.0% even though the training set of TableLLM contains 18k cases from WikiTQ and 28k cases from WikiSQL. And our TrixInstruct only contains 6k training instances in total. It demonstrates the effectiveness of finetuning with Plan-then-Reason framework. ProTrix-Coder also surpasses TableLLM on fact verification tasks by 13.8% on average.

### v3

@@ Abstract
Tables play a crucial role in conveying information of various domains. We propose a novel Plan-then-Reason framework to answer different types of user queries over tables with sentence context. Our framework first plans the reasoning paths over both tabular and sentence context, then assigns each step to program-based or textual reasoning to reach the final answer. This framework enhances the table reasoning abilities for both in-context learning and finetuning methods. GPT-3.5-Turbo following our Plan-then-Reason framework surpasses other prompting baselines that do not employ self-consistency while using fewer API calls and in-context demonstrations. We also construct an instruction tuning set TrixInstruct to evaluate the effectiveness of finetuning with this framework. We present a ProTrix model family by finetuning models on TrixInstruct. Our experiments show that the ProTrix family can generalize to diverse unseen tabular tasks with only 6k training instances. We further demonstrate that ProTrix can generate accurate and faithful explanations to answer complex free-form questions. Our work underscores the importance of the planning and reasoning abilities towards a model over tabular tasks with generalizability and interpretability. We open-source our dataset and models at https://github.com/WilliamZR/ProTrix.
@@ 1 Introduction
Tables, serve as a fundamental tool for organizing and presenting information across various domains. Whether in business reports, or scientific publications, tables are commonly employed to represent complex data effectively. Despite their widespread utility, the process of human beings answering questions involving tables appears to be time-consuming, given the often substantial amount of content involved. Recognizing this challenge, there arises a need to leverage Large Language Models (LLMs) to understand and respond to user queries automatically.
Figure 1: Demonstrations of user queries to a table in Wikipedia. Some of the sentences with hyperlinks to the table are presented as sentence context.
@@
The last query also requires the model to recall commonsense knowledge to decide which contextual information can be used as evidence to judge if Kenya is successful at the 2001 Goodwill Games. Subsequently, the model must generate explanations to arrive at certain conclusions. The first two queries mainly require the model to fill the information gap in the query with a short-form answer while the third query seeks for information on a more general concept. The queries underscore the importance of planning and reasoning abilities to connect the query with the actual information in the context and generate faithful and accurate explanations for conclusions.
@@
There are generally two ways to enhance a model’s reasoning ability. One is textual reasoning which prompts the model to answer questions step-by-step Wei et al. (2022). The other one is program-based reasoning, prompting the model to write code to answer the questions Chen et al. (2022). However, both reasoning methods have their own disadvantages. The textual reasoning method such as Chain-of-Thought Wei et al. (2022) can be used to enhance the tabular reasoning ability but often lacks precision in tabular operations such as sorting, counting and filtering as shown in the right example in Figure 2, and may not generalize well to large tables Chen (2023). The program-based reasoning method can reason with high precision with SQL or Python code Chen et al. (2022). The left one in Figure 2 queries the color comparison between steaks with different cooking methods which is program-unsolvable by SQL and Python. Therefore, it requires the model to adaptively plan the usage of the program-based and textual reasoning to make the best of both methods.
Previous works incorporating LLMs’ dynamic planning abilities for table reasoning Yao et al. (2023); Zhang et al. (2023b); Wang et al. (2024) primarily focus on selecting local actions during the reasoning process, rather than globally planning the entire reasoning strategy beforehand. In this paper, we propose a novel Plan-then-Reason framework, which decomposes the question and adapts the reasoning chain based on the table and sentence context in advance, preventing the model from becoming trapped in localized reasoning steps Zhang et al. (2023c). Instead of fixed reasoning patterns Wei et al. (2022); Chen et al. (2022); Wang et al. (2023); Nahid and Rafiei (2024), our approach offers greater flexibility by adaptively assigning reasoning sub-tasks to either SQL-based or text-based methods, depending on the nature of the decomposed sub-questions. Experiments demonstrate that our Plan-then-Reason framework generates more accurate answers with fewer API calls, surpassing existing prompting methods while requiring fewer in-context examples.
@@
Meanwhile, few finetuning methods are designed to enhance both planning and reasoning abilities of open-source models while they are crucial for building tabular models with generalizability and interpretability. Various pre-trained models are proposed for tabular tasks Yin et al. (2020); Wang et al. (2021); Iida et al. (2021); Deng et al. (2022); Yang et al. (2022); Jiang et al. (2022); Liu et al. (2021). But they are limited to specific query types and could not generalize well to unseen tasks. Models finetuned with respect to general tabular querying tasks Xie et al. (2022); Liu et al. (2023a); Zhang et al. (2023a); Zhuang et al. (2024) are expected to generate answers directly, which inevitably lacks interpretability. The reasoning method for TableLLM Zhang et al. (2024) is primarily determined by the size of tables, rather than being planned based on the query and context.
@@
Recent base models are pre-trained with a large amount of corpora thus obtaining intrinsic commonsense knowledge Touvron et al. (2023); Roziere et al. (2023). These models suit as the backbones for our framework that can fill the gap between queries and tables, understand general concepts, and plan the reasoning paths over table and sentence context. We construct an instruction tuning dataset TrixInstruct based on benchmarks with queries that are program-unsolvable or need combining information from table and sentence context. We finetune Llama-2-7B Touvron et al. (2023) and CodeLlama-7B Roziere et al. (2023) with TrixInstruct. The resulting ProTrix [^1] Protrix originally means a chemical reactor for small-scale production with compatibility and process control. model family is designed to Plan and Reason On TabulaR tasks with integration of code executIon and teXtual reasoning. Our experiments show that models trained with our Plan-then-Reason framework can generalize to unseen tabular tasks in different domains with only a handful of training examples and give accurate and faithful explanations even for complex how and why questions.
@@
$\bullet$ We propose a Plan-then-Reason framework to answer user queries on tabular tasks with sentence context. The framework first plans the reasoning pathways by ingesting the query and the context, and assigns each step to textual and program-based reasoning to arrive at the final answer. We experiment with GPT-3.5-turbo to evaluate the effectiveness of this framework and find it surpasses existing methods that do not employ self-consistency.
@@ 2.1 Problem Formulation
This study centers on tabular tasks with sentence context. Each instance is structured as $(Q,T,S,A)$, where $Q$ represents users’ query. $T$ denotes a singular table, while $S$ denotes the sentence context. The sentence context is usually passages linked to the table or retrieved from a knowledge base. Finally $A$ stands for the answer, which could be short-form for answering questions like hom many… or is this true or false…. For how and why questions, the answers are generally one or more sentences, defined as free-form answers.
Figure 3: Illustration of our framework. The figure shows the process of our framework to answer a question. The framework first performs strategical planning, decomposing questions into reasoning chains to be solved by either table or sentence context, then performs program-based or textual reasoning to reach the answers.
@@ 2.2 Plan-then-Reason
We propose a Plan-then-Reason framework to answer queries over tables and texts. The framework first ingests the query and the context by recalling common knowledge and general concepts. Then it begins to design the model’s reasoning pathway, planning the utilization of program-based and textual reasoning to arrive at conclusions.
@@
Then the model adaptively plans the reasoning path with both program-based and textual reasoning to address the limitations of each individual reasoning method. For the first query, the model plans to use SQL to extract relevant cells from the table and make comparisons of concepts through textual reasoning. For the second query, the model decomposes the task into a multi-hop reasoning chain. It uses SQL to extract the silver medalist from the table and uses sentence context to verify his Olympic record.
@@
The reasoning phase initiates with program-based reasoning, writing SQL queries to extract relevant cells or perform operations such as counting and sorting. After running SQL on a code interpreter, the results are fed back into the model’s ongoing reasoning process. Subsequently, during textual reasoning, the model selects relevant sentences from noisy context to complement the table context. Reuben Kosgei… youngest ever winner of an Olympic gold medal suggests that he has won an Olympic gold medal in his career. Finally, the model summarizes insights from program-based and textual reasoning to reach the answer.
@@ 2.3 In-Context Learning
One intuitive way to evaluate the effectiveness of our Plan-then-Reason work is through in-context learning. We prompt LLMs to follow the planning and reasoning pattern in Figure 3. We experiment with GPT-3.5-turbo on WikiTQ Pasupat and Liang (2015), FEVEROUS Aly et al. (2021) and TabFact Chen et al. (2019) following the prompts shown in Table 1112 and 13 in Appendix.
@@ 2.4 Instruction Tuning
Based on the analysis in §1, we highlight the abilities our Plan-then-Reason framework possesses towards tabular tasks with sentence context. (1) Understanding user’s query: use parametric knowledge of commonsense insights and general concepts to analyze the relationship between the query and the context; (2) Adaptive planning: decompose a query into sub-questions and plan to answer these sub-questions with different types of context or design multi-hop reasoning paths for the query, and (3) Blending program-based and textual reasoning: synthesize the strength of each method to maximize reasoning performance.
To train our model with such abilities, we construct an instruction tuning set TrixInstruct based on two benchmarks i.e., WikiTQ Pasupat and Liang (2015) and FEVEROUS Aly et al. (2021). WikiTQ involves table question-answering based on a single table, requiring multi-step reasoning and performing diverse data operations such as comparison, aggregation, and arithmetic computation. WikiTQ also contains cases that can not be solved by SQL programming solely Shi et al. (2020) which require extra textual reasoning as the left example in Figure 2. Finetuning with such cases equips our models with the ability to plan between textual or program-based reasoning.
@@
For each task, we sample 4,000 instances from the training datasets. We employ GPT-4 Achiam et al. (2023) to generate responses according to the Plan-then-Reason framework following the prompts in Table 11 and 12 in Appendix. We filter out instances that GPT-4 predicts answers inconsistent with the original annotations. This results in a training set comprising 3,157 cases from FEVEROUS and 2,866 cases from WikiTQ. We train our models based on Llama-2-7B Touvron et al. (2023) and CodeLlama-7B Roziere et al. (2023) resulting in two models ProTrix and ProTrix-Coder.
@@ 3.1 Benchmarks for Evaluation
WikiSQL and WikiTQ are question answering benchmarks on tables from Wikipedia without sentence context Zhong et al. (2017); Pasupat and Liang (2015). HybridQA Chen et al. (2020) requires models to answer questions based on both tables and sentences. We use retrieved sentences, admittedly noisy, from Chen et al. (2020) as the sentence context. TATQA Zhu et al. (2021) focuses on tables with sentence context from financial reports.
@@
We use the same method in §2.4 to construct the evaluation dataset for FEVEROUS Aly et al. (2021). TabFact Chen et al. (2019) verifies claims based on tables from Wikipedia. SciTAB Lu et al. (2023) focuses on tables from scientific papers. This benchmark requires compositional reasoning and commonsense knowledge.
@@
FetaQA contains what questions with multiple answers and how and why questions that require models to generate explanations Nan et al. (2022). The original FetaQA dataset has highlighted relevant cells, we turn to a more challenging and realistic scenario where the highlighted cells are not provided as input and the model will answer the question directly based on the complete table context. Since our model is only finetuned on short-form answer tasks, FetaQA can be utilized to further evaluate the interpretability and generalizability of our models.
@@ 3.2 Short-form Answer Tasks
We choose the following baselines: (1) Closed-source model: We use the end-to-end QA performance of GPT-3.5-turbo and GPT-4 as baselines and compare our results on GPT-3.5-turbo with Binder Cheng et al. (2022b), ReAcTable Zhang et al. (2023b), StructGPT Jiang et al. (2023), Dater Ye et al. (2023), Chain-of-Table Wang et al. (2024) and TabSQLify Nahid and Rafiei (2024). (2) Finetuned SOTA: Please refer to Appendix A for details of finetuned SOTA methods. (3) 7B parameter models: We first compare our models with the zero-shot performance of base models, Llama-2-7B Touvron et al. (2023) and CodeLlama 7B Roziere et al. (2023). Then we compare with TableLlama Zhang et al. (2023a) and TableLLM Zhang et al. (2024), which are the most similar baselines to our models and share the same base models [^2] We do not use StructLM Zhuang et al. (2024) as baseline since it is finetuned on most of our evaluation benchmarks while we want to compare with existing models in terms of generalizability.. TableLlama is a generalist model for end-to-end QA. TableLLM originally chooses textual or program-based reasoning for different benchmarks. To test the generalizability for different queries, we choose the reasoning method of TableLLM by table size. We use textual reasoning for tables less than 500 tokens and program-based reasoning for larger tables. We provide a detailed comparison with other closed-source models (Codex, PaLM2) and open-source model (Llama-3-8B) in Table 10 in Appendix.
@@
For fact verification, we match the predicted veracity label in the concluding sentence, and use the evaluator from Cheng et al. (2022b) to evaluate the performance of Plan-then-Reason with GPT-3.5-turbo on WikiTQ. For 7B parameter models, we find that the heuristic matching evaluation for question-answering tasks would lead to overestimation or underestimation since our model is not finetuned to follow the grammar of gold answers in each dataset. Similar to Zhang et al. (2024), we deploy Llama-2-70B-chat Touvron et al. (2023) to predict if the concluding sentence answers the question correctly using the prompt in Table 14 in Appendix. Our evaluation results using LLM are checked manually and they align better with human evaluation of the concluding sentence. We report three-class F1 score for SciTAB and accuracy for other datasets.
@@
The result with GPT-3.5-turbo [^3] We use GPT-3.5-turo-16k-0613 for experiment. We only experiment with WikiTQ, TabFact and FEVEROUS due to limited budgets. in Table 1 suggests that Plan-then-Reason framework is effective for answering different types of queries over tables. Our method surpasses all the baselines without self-consistency on WikiTQ. It also achieves higher accuracy than Dater and Binder, which use 20 responses and 50 responses for self-consistency, respectively. Our method still falls behind Mix SC which predicts answers based on 5 responses with textual reasoning and 5 with program-based reasoning. But the self-consistency mechanism leads to a much larger computing cost. Our framework also surpasses existing baselines on TabFact except for StructGPT, demonstrating its effectiveness for fact verification tasks. Plan-then-Reason can also generalize to tables with sentence context. It increases the accuracy on FEVEROUS by 4.8%.
@@
The experimental results with Llama-2-7B and CodeLlama-7B in Table 1 show that our finetuned models generalize to diverse tabular tasks with only 6k training instances. Compared with the backbone model Llama-2-7B, the performance gain of ProTrix on in-domain benchmarks is 34.8% on WikiTQ and 28.7% on FEVEROUS. And the performance gain on out-of-domain benchmarks is 21.5% on average. Comparing the out-of-domain performance with TableLlama, ProTrix surpasses TableLlama by 25.7% on WikiSQL, 15.8% on SciTAB, 9.6% on HybridQA and 11.8% on TATQA. The overall performance gain on out-of-domain benchmarks demonstrates the planning and reasoning abilities obtained from TrixInstruct is not restricted to in-domain tasks. Our finetuned model ProTrix adaptatively generalizes to queries with different input and output configurations and can even be applied to specific domains such as science and finance.
ProTrix-Coder still falls behind TableLLM on question answering task with sentence context since TableLLM is finetuned with 8k cases from TATQA. We are surprised to find that ProTrix-Coder surpasses TableLLM on WikiTQ and WikiSQL by 4.9% and 7.0%, respectively, even though the training set of TableLLM contains 18k cases from WikiTQ and 28k cases from WikiSQL. Our TrixInstruct only contains 6k training instances in total. These results indicate the effectiveness of finetuning with our Plan-then-Reason framework. ProTrix-Coder also surpasses TableLLM on fact verification tasks by 13.8% on average.